# envname-P envname#1 BOAD: Discovering Hierarchical Software Engineering Agents via Bandit Optimization

### v1

@@ Abstract
Large language models (LLMs) have shown strong reasoning and coding capabilities, yet they struggle to generalize to real-world software engineering (SWE) problems that are long-horizon and out of distribution. Existing systems often rely on a single agent to handle the entire workflow—interpreting issues, navigating large codebases, and implementing fixes—within one reasoning chain. Such monolithic designs force the model to retain irrelevant context, leading to spurious correlations and poor generalization. Motivated by how human engineers decompose complex problems, we propose structuring SWE agents as orchestrators coordinating specialized sub-agents for sub-tasks such as localization, editing, and validation. The challenge lies in discovering effective hierarchies automatically: as the number of sub-agents grows, the search space becomes combinatorial, and it is difficult to attribute credit to individual sub-agents within a team. We address these challenges by formulating hierarchy discovery as a multi-armed bandit (MAB) problem, where each arm represents a candidate sub-agent and the reward measures its helpfulness when collaborating with others. This framework, termed Bandit Optimization for Agent Design (BOAD), enables efficient exploration of sub-agent designs under limited evaluation budgets. On SWE-bench-Verified, BOAD outperforms single-agent and manually designed multi-agent systems. On SWE-bench-Live, featuring more recent and out-of-distribution issues, our 36B system ranks second on the leaderboard at the time of evaluation, surpassing larger models such as GPT-4 and Claude. These results demonstrate that automatically discovered hierarchical multi-agent systems significantly improve generalization on challenging long-horizon SWE tasks. Code is available at https://github.com/iamxjy/BOAD-SWE-Agent.
@@ 1 Introduction
We evaluate our BOAD on SWE-bench-Verified [25], a benchmark for software engineering tasks grounded in real GitHub issues. On SWE-bench-Verified, our method consistently outperforms single-agent systems and manually designed multi-agent systems. On SWE-bench-Live [30], which includes more recently collected issues and presents out-of-distribution challenges, our system achieves 3rd place on the leaderboard using a 36B model—outperforming larger-scale systems based on Claude and GPT-4. To our best knowledge, our work is the first to show generalization improvements using automatically discovered hierarchical multi-agent systems on challenging long-horizon interactive tasks like SWE-bench.
@@ 3 Preliminaries
Semi-Markov Decision Process (SMDP) framework [39] is widely used to mitigate long-horizon sparse reward problems. Instead of issuing primitive actions , the orchestrator selects a temporally extended action (option) . Each option corresponds to a sub-agent that executes a sequence of actions until termination, after which control returns at , where denotes the duration of an sub-agent. This reduces decision frequency and simplifies planning.
@@ 5.2 Main results
We also implement an evolutionary search for a multi-agent system adapted from Automated Design of Agentic Systems [23], as a baseline to compare BOAD against. Implementation details for the evolutionary search are provided in Appendix A.4.1. When using the same number of evaluation instances (i.e number of SWE-Bench patches generated), we find that the sub-agents generated by the evolutionary search achieve worse performance (17.0% vs 20.0% on SWE-Bench-Live) than BOAD. Additionally, for the same number of iterations, the cost of Claude API calls is more than double than that of BOAD due to the need of generating many sub-agents at each iteration, whereas BOAD reuses subagents across iterations.
@@ 5.4 Qualitative Analysis of Single- vs Multi-Agent Outcomes
Multi-site fixes and coverage (multi-agent advantage). When the fix required edits at multiple call sites or modules, the single agent often either over-edited unrelated regions or missed one or more necessary locations. The hierarchical system mitigated both omission and extraneous edits by delegating to a sub-agent for localization, which did a thorough analysis of the repository before making any targetted modifications.
@@
Error propagation from unvalidated sub-agent outputs (multi-agent failure mode). In a minority of cases, the multi-agent system failed while the single agent succeeded. Inspecting these outputs, we found that erroneous sub-agent outputs (e.g., incomplete span identification, misinterpretation of the issue) were accepted as ground truth by the orchestrator, leading subsequent steps astray. Because there is no intermediate validation or self-checking, the orchestrator has limited ability to recover from such upstream.
@@ A.1 Prompt templates
A.1.1 Prompt template for refining subagent during warmup stage
@@ A.1.3 Custom orchestrator plan prompt
A.1.4 Prompt template for checking if a subagent was helpful in a given instance
@@ A.1.5 Orchestrator-only Prompt
A.2 Subagents
A.2.1 BOAD Top 2 Discovered Subagent Configurations for Seed-OSS-36B-Instruct
A.2.2 All BOAD Discovered Subagents
@@
| Subagent | Generated Iteration | Helpfulness | |
@@
A.2.3 All Evolution Discovered Subagents
Here we report all bundles discovered over 20 iterations of BOAD (one for each iteration). For each sub-agent, we report its iteration number, the generated subagents, and the bundle’s success rate. Note that sub-agents with the same name across different iterations may have different system/instance templates because sub-agents are not reused. On evaluation, we choose the bundle with the highest success rate (latest bundle if tied).
@@ A.2.3 All Evolution Discovered Subagents
| Iteration | Subagents | Success Rate |
@@
A.2.4 Manually Designed Subagent Configurations
@@ A.4.1 Implementation Details for Evolutionary Baseline
For comparison with BOAD, we implemented an evolutionary multi-agent design baseline where the orchestrator and three sub-agent prompts are bundled together and treated as a single evolutionary individual. We use three sub-agents paired with an orchestrator because BOAD chooses three sub-agents when optimizing the sub-agents in the experiments (Section 5.2). At each iteration of the evolution, the system generates three new subagents sequentially, given the previous round subagent configurations and their measured helpfulness and success rates (using the same prompts as BOAD, provided in A.1.2). Next, both the sub-agent warm-up and the orchestrator construction follow the same procedures used in BOAD. The generated orchestrator and subagents are then run on the same design set to get the helpfulness and success rate for the next iteration.
@@ A.4.2 API Cost Comparison with BOAD
The evolutionary baseline also requires Claude calls—both to propose new orchestrator/sub-agent prompts and to perform LLM-as-judge scoring. Under the same setup, each evolution iteration costs $2.33, whereas each BOAD iteration costs $0.96, indicating that evolution is substantially more expensive to run. Evolution incurs higher cost because it mutate the bundle of orchestrator and sub-agents at each iteration.

### v2

@@ Abstract
Large language models (LLMs) have shown strong reasoning and coding capabilities, yet they struggle to generalize to real-world software engineering (SWE) problems that are long-horizon and out-of-distribution. Existing systems often rely on a single agent to handle the entire workflow—interpreting issues, navigating large codebases, and implementing fixes—within one reasoning chain. Such monolithic designs force the model to retain irrelevant context, leading to spurious correlations and poor generalization. Motivated by how human engineers decompose complex problems, we propose structuring SWE agents as orchestrators coordinating specialized sub-agents for sub-tasks such as localization, editing, and validation. The challenge lies in discovering effective hierarchies automatically: as the number of sub-agents grows, the search space becomes combinatorial, and it is difficult to attribute credit to individual sub-agents within a team. We address these challenges by formulating hierarchy discovery as a multi-armed bandit (MAB) problem, where each arm represents a candidate sub-agent and the reward measures its helpfulness when collaborating with others. This framework, termed Bandit Optimization for Agent Design (BOAD), enables efficient exploration of sub-agent designs under limited evaluation budgets. On SWE-bench-Verified, BOAD outperforms single-agent and manually designed multi-agent systems. On SWE-bench-Live, featuring more recent and out-of-distribution issues, our 36B system ranks second on the leaderboard at the time of evaluation, surpassing larger models such as GPT-4 and Claude. These results demonstrate that automatically discovered hierarchical multi-agent systems significantly improve generalization on challenging long-horizon SWE tasks. Code is available at https://github.com/iamxjy/BOAD-SWE-Agent.
@@ 1 Introduction
We evaluate our BOAD on SWE-bench-Verified [25], a benchmark for software engineering tasks grounded in real GitHub issues. On SWE-bench-Verified, our method consistently outperforms single-agent systems and manually designed multi-agent systems. On SWE-bench-Live [30], which includes more recently collected issues and presents out-of-distribution challenges, our system ranked second on the leaderboard at the time of evaluation using a 36B model—outperforming larger-scale systems based on Claude and GPT-4. To our best knowledge, our work is the first to show generalization improvements using automatically discovered hierarchical multi-agent systems on challenging long-horizon interactive tasks like SWE-bench.
@@ 3 Preliminaries
The Semi-Markov Decision Process (SMDP) framework [39] is widely used to mitigate long-horizon sparse reward problems. Instead of issuing primitive actions , the orchestrator selects a temporally extended action (option) . Each option corresponds to a sub-agent that executes a sequence of actions until termination, after which control returns at , where denotes the duration of a sub-agent. This reduces decision frequency and simplifies planning.
@@ 5.2 Main results
We also implement an evolutionary search for a multi-agent system adapted from Automated Design of Agentic Systems [23], as a baseline to compare BOAD against. Implementation details for the evolutionary search are provided in Appendix A.4.1. When using the same number of evaluation instances (i.e., number of SWE-Bench patches generated), we find that the sub-agents generated by the evolutionary search achieve worse performance (17.0% vs 20.0% on SWE-Bench-Live) than BOAD. Additionally, for the same number of iterations, the cost of Claude API calls is more than double that of BOAD due to the need of generating many sub-agents at each iteration, whereas BOAD reuses subagents across iterations.
@@ 5.4 Qualitative Analysis of Single- vs Multi-Agent Outcomes
Multi-site fixes and coverage (multi-agent advantage). When the fix required edits at multiple call sites or modules, the single agent often either over-edited unrelated regions or missed one or more necessary locations. The hierarchical system mitigated both omission and extraneous edits by delegating to a sub-agent for localization, which did a thorough analysis of the repository before making any targeted modifications.
@@
Error propagation from unvalidated sub-agent outputs (multi-agent failure mode). In a minority of cases, the multi-agent system failed while the single agent succeeded. Inspecting these outputs, we found that erroneous sub-agent outputs (e.g., incomplete span identification, misinterpretation of the issue) were accepted as ground truth by the orchestrator, leading subsequent steps astray. Because there is no intermediate validation or self-checking, the orchestrator has limited ability to recover from such upstream errors.
@@ A.1 Prompt templates
A.1.1 Prompt template for refining sub-agent during warmup stage
@@ A.1.3 Custom orchestrator plan prompt
A.1.4 Prompt template for checking if a sub-agent was helpful in a given instance
@@ A.1.5 Orchestrator-only Prompt
A.2 Sub-agents
A.2.1 BOAD Top 2 Discovered Sub-agent Configurations for Seed-OSS-36B-Instruct
A.2.2 All BOAD Discovered Sub-agents
@@
| Sub-agent | Generated Iteration | Helpfulness | |
@@
A.2.3 All Evolution Discovered Sub-agents
Here we report all bundles discovered over 20 iterations of BOAD (one for each iteration). For each sub-agent, we report its iteration number, the generated sub-agents, and the bundle’s success rate. Note that sub-agents with the same name across different iterations may have different system/instance templates because sub-agents are not reused. On evaluation, we choose the bundle with the highest success rate (latest bundle if tied).
@@ A.2.3 All Evolution Discovered Sub-agents
| Iteration | Sub-agents | Success Rate |
@@
A.2.4 Manually Designed Sub-agent Configurations
@@ A.4.1 Implementation Details for Evolutionary Baseline
For comparison with BOAD, we implemented an evolutionary multi-agent design baseline where the orchestrator and three sub-agent prompts are bundled together and treated as a single evolutionary individual. We use three sub-agents paired with an orchestrator because BOAD chooses three sub-agents when optimizing the sub-agents in the experiments (Section 5.2). At each iteration of the evolution, the system generates three new sub-agents sequentially, given the previous round sub-agent configurations and their measured helpfulness and success rates (using the same prompts as BOAD, provided in A.1.2). Next, both the sub-agent warm-up and the orchestrator construction follow the same procedures used in BOAD. The generated orchestrator and sub-agents are then run on the same design set to get the helpfulness and success rate for the next iteration.
@@ A.4.2 API Cost Comparison with BOAD
The evolutionary baseline also requires Claude calls—both to propose new orchestrator/sub-agent prompts and to perform LLM-as-judge scoring. Under the same setup, each evolution iteration costs $2.33, whereas each BOAD iteration costs $0.96, indicating that evolution is substantially more expensive to run. Evolution incurs higher cost because it mutates the bundle of orchestrator and sub-agents at each iteration.